There are 10 students in a class. How many ways can they be arranged in a...

  1. There are 10 students in a class. How many ways can they be arranged in a line if two friends always want to stay together?

    I.             100

    II.          3,628,800

    III.       725,760

    IV.       45

    V.          90

    VI.       10

    I
    II
    III
    IV
    VI
    VI

  2. There are 10 students in a class. If each student shakes hand with every other student, how many possible shake-hands have taken place?

    I.                   100

    II.                3,628,800

    III.             725,760

    IV.             45

    V.                90

    VI.             10

    I
    II
    III
    IV
    V
    VI
Answer #1

Two friends want to sit together. So, we can consider them as one person
For 9 members to sit in a line, it is 9!
II. 3,628,800

Hand shakes for 10 persons
Each one shakes hands with other 9: n(n-1)
For two persons it is considered twice. So n(n-1)/2
n(n-1)/2 = 10(9)/2 = 5*9 = 45
IV. 45